# Condition monitoring of wind turbine blades via learning-based methods

### v1

@@ Abstract
journal: Renewable Energy
@@ 1 Introduction
As the number of wind turbines continues to rise, there is a growing need to identify new suitable locations for new wind farms. The limited availability of optimal sites is reflected in the trend of constructing taller turbines with bigger blades for making the investment economically viable in less optimal areas. The design of bigger blades requires different solutions, with the consequence of dealing with different fault scenarios that have higher economical impact. Even with the current blade designs state, about $14.9\%$ of the faults on WT are blade failures [2], caused by both endogenous and exogenous factors [3, 4, 5]. Furthermore, blades are the structural element of the turbine most affected by damage, due to their high complexity and prolonged exposition to strong dynamical loads [3]. As a consequence, interest arose in the investigation of condition monitoring and fault detection strategies to be able to detect faults preemptively before they become more severe and thus considerably more costly.
Model-based Fault Detection (FD) strategies are the most traditional schemes used in wind turbines. Starting form a linearised model of the system, residuals are commonly generated starting from Kalman filters and the collected data [6]. In [7] and [8] a closed-loop scheme that describes the WT model is provided, which was then used for residual generation using Kalman Filters. In the latter, wind speed was taken as a reliable measurement for the design of the Kalman filter. However, it is generally difficult to obtain accurate local wind speed measurements [9]. This issue was addressed in [10], where the authors estimated the wind speed via estimating the aerodynamic torque.
@@
Neural network have also been explored because of their capability of dealing with underlying nonlinear functions. Autoencoders were applied in [17], to detect faults occurring on the gearbox and generator. Other works have continued on this path, training the autoencoders to recognise different faults happening on the WT, such as [18], and [19]. An AE model trained with Long Short-Term Memory layers was proposed in [20] to recognise when a blade cracks.
Although data-driven methods for fault detection in wind turbines has been studied in the recent literature, dedicated studies on condition monitoring of the blades has been sparsely reported. This paper explicitly focuses on the design and application of learning-based fault diagnosis in WT. The main contributions of this paper are
@@ 2 System description and problem statement
    Pitch angles of each blade (3 measurements)
According to the wind speed, the power output and the rotor speed, 5 different operating regions are defined, where the turbine is operated with different control strategies. Such signals, together with the consequent blade pitch angle are normalized and presented in Figure 1, highlighting the different operating regions at which the turbine is operating at.
@@
A long recording with the turbine not subject to fault is utilized for fitting the learning-based models. Such data are normalized and split into training, validation and testing for designing and verifying the design under nominal conditions. Some faulty data are available for the same turbine subject to different blade faults. These data were not accounted for in the design stage and are used to test the models under real-life faulty scenarios.
@@ 2.1 Failure Modes
The fault scenario that were considered in the investigation can be divided in three main categories:
@@
    Two common phenomenon that might affect the blade are icing, caused by the formation of ice because of low temperatures, and the formation of cracks along the surface of the blades. Such events could entail a rather sudden change in the moment of inertia of the blade or its aerodynamic properties, hence showing a sudden jump in the measurement of the root bending moments, as stated in [21]. Hence, they were simulated on the available data from the healthy turbine by adding an offset to the original bending moment measurements. A bias in the measurement could also be caused by a fault occurring in the sensor itself.
@@
    Another common fault is represented by the case of a sensor measurements being stuck because of its incorrect setup on the blade, or by defects in the sensors itself. This kind of fault was simulated by changing the measured bending moment signal to become constant after the fault occured.
@@
    Finally, it was decided to use real-life data collected from a blade undergoing a failure. For this reason, Vestas Wind Systems A/S provided some data collected with the turbine experiencing a fault.
@@ 3 Methods and theoretical overview
Once the model is trained, the testing Stage is performed online: new data are collected from the wind turbine, into an unknown state, and then reconstructed by the trained model. The fault detection algorithm is obtained by computing a suitable statistic to compare the data reconstructed by the trained model and the real data measured by the sensors (usually by means of Square Prediction Error or Mean Average Error), and by comparing the result with the fault-free case obtained in the training stage by means of a GLR test. If the two statistics are similar, then the GLRT assumes that the new data are from a healthy plant, while if they are showing strong differences, it detects a fault. Two approaches are proposed (PCA and Autoencoders), differing both in the model structure and in the training method.
@@ 3.1 Introduction to PCA
Principal Component Analysis (PCA) is an unsupervised, linear method for dimensional reduction in large multivariate datasets. It transforms a normalized data matrix $\boldsymbol{X}\in\mathbb{R}^{n\times m}$ (with $n$ samples and $m$ features) into a lower-dimensional space by applying singular value decomposition (SVD) to its covariance matrix:
@@
$SPE(k)=\sum_{i=1}^{m\omega}\left(\hat{X}_{d}^{(i)}(k)-X_{d}^{(i)}(k)\right)^{2}=\boldsymbol{X_{d}}(k)(\boldsymbol{I}-\boldsymbol{\hat{P}\hat{P}}^{T})(\boldsymbol{I}-\boldsymbol{\hat{P}\hat{P}}^{T})^{T}\boldsymbol{X_{d}}(k)^{T}$ (4)
@@ 5 Experimental results
Both the PCA and the AE models are tested for the faulty scenarios presented in Section 2.1. While defining the experiment for the third category is straightforward, thanks to the availability of faulty data, the first two categories require the corruption of (unused) healthy data to mimic the previously described faulty scenarios. An overview of the experiment setting for each faulty scenario is presented in Table 2, where $M_{og}(t)$ represents the blade root bending moment measured on the original healthy turbine at time $t$, while $t_{f}$ is the time instant where the fault occurs.
@@
Real-life scenario 1: catastrophic fault Sudden jump in the flap bending moment signal
Real-life scenario 2: slower-forming Exponential increase in the flap bending moment
Starting from the real-life scenario, both the reconstruction error and the result of applying the GLR are presented in Figure 7. The reconstruction error changes in both cases and the filtered version allows to detect the fault by the direct use of a threshold. The GLR proves its effectiveness in improving signal separation so that the rate of false alarms is lowered.
@@
The same models are applied to the (new) healthy data corrupted with sensor faults. Because of the better detection performance obtained with the GLR, the direct threshold methods results are omitted. The results are presented in Figure 8. The capability of the AE of learning the nonlinear features of the model is reflected in a better performance over all the possible faults. In particular, the PCA is in general not able to detect the fault associated with the Edge-bending-moment sensor, while it is still able to detect faults in the flap-bending moment sensor by leveraging the correct model according to the operating region the wind turbine is working at. Even under such conditions, the AE is better performing by ensuring strong detectability and no false alarms in the test that have been performed.
@@ 6 Conclusions
This paper presented a condition monitoring framework for detecting faults in wind turbine blades based on learning based-models, trained to reconstruct a set of measurements under healthy conditions and showing discrepancies otherwise.
Starting with pre-processed real-life data from a healthy wind turbine, both a linear model based on PCA and a nonlinear model based on autoencoders were trained to reconstruct the input data. The trained models were then tested on faulty data. Although both PCA and autoencoder-based models demonstrated their capability in detecting faults in wind turbine blades, the autoencoder-based approach proved to be the most effective thanks to its ability of dealing with all the tested fault scenarios and by leveraging a single model for all the operating regions.
Future work involves optimizing the algorithm for real-time implementation and testing different model architectures to improve detection performance.

### v2

@@ Abstract
journal: Engineering Applications of Artificial Intelligence
@@ 1 Introduction
As the number of wind turbines continues to rise, there is a growing need to identify new suitable locations for new wind farms. The limited availability of optimal sites is reflected in the trend of constructing taller turbines with bigger blades for making the investment economically viable in less optimal areas. The design of bigger blades requires different solutions, with the consequence of dealing with different fault scenarios that have higher economical impact. Even with the current blade designs state, about $14.9\%$ of the faults on WT are blade failures [2], caused by both endogenous and exogenous factors [3, 4, 5]. Furthermore, blades are the structural element of the turbine most affected by damage, due to their high complexity and prolonged exposition to strong dynamical loads [3]. As a consequence, interest arose in the investigation of condition monitoring and fault detection strategies to be able to detect faults preemptively before they become more severe and thus considerably more costly. A comprehensive review of the research methods is presented in [6, 7].
Model-based Fault Detection (FD) strategies are the most traditional schemes used in wind turbines and their subsystems [8]. Starting form a linearised model of the system, residuals are commonly generated starting from Kalman filters and the collected data [9]. In [10] and [11] a closed-loop scheme that describes the WT model is provided, which was then used for residual generation using Kalman Filters. In the latter, wind speed was taken as a reliable measurement for the design of the Kalman filter. However, it is generally difficult to obtain accurate local wind speed measurements [12]. This issue was addressed in [13], where the authors estimated the wind speed via estimating the aerodynamic torque.
@@
Neural network have also been explored because of their capability of dealing with underlying nonlinear functions. Autoencoders (AE) were applied in [20], to detect faults occurring on the gearbox and generator. Other works have continued on this path, training the autoencoders to recognise different faults happening on the WT, such as [21], and [22]. An AE model trained with Long Short-Term Memory (LSTM) layers was proposed in [23] to recognise when a blade cracks.
Although data-driven methods for fault detection in wind turbines has been studied in the recent literature, dedicated studies on condition monitoring of the blades has been sparsely reported. This paper explicitly focuses on the design and application of learning-based fault diagnosis scheme for detecting blade faults in wind turbines. The main contributions of this paper are
@@ 2 System description and problem statement
    Pitch angles of each blade (3 measurements).
Such signals have been chosen because of their relevance, both in terms of being able to display a change under blade fault and for determining the wind turbine operating point correctly. In fact, according to the wind speed, the power output and the rotor speed, 5 different operating regions are defined, where the turbine is operated with different control strategies, and thus is in general characterized by a different underline dynamics. Such signals, together with the consequent blade pitch angle are normalized and presented in Figure 1, highlighting the different operating regions at which the turbine is operating at.
@@
A long recording with the turbine not subject to fault will be utilized for fitting the learning-based models. Such data are normalized and split into training, validation and testing for designing and verifying the design under nominal conditions. Some faulty data are available for the same turbine subject to different blade faults. These data were not accounted for in the design stage and are used to test the models under real-life faulty scenarios.
@@ 2.1 Failure Modes
The fault scenarios that were considered in the investigation can be divided in three main categories:
@@
    Two common phenomenon that might affect the blade are icing, caused by the formation of ice because of low temperatures, as stated in [3], and the formation of cracks along the surface of the blades. Such events could entail a rather sudden change in the moment of inertia of the blade or its aerodynamic properties, hence showing a sudden jump in the measurement of the root bending moments, as stated in [24]. Hence, they were simulated on the available data from the healthy turbine by adding an offset to the original bending moment measurements. A bias in the measurement could also be caused by a fault occurring in the sensor itself. The fault described by the biased root bending moment measurement is hence representing both a faulty operation derived from normal physical phenomenon and a the corruption of a critical sensor for the turbine.
@@
    Another common fault is represented by the case of a sensor measurement being stuck because of its incorrect setup on the blade, or by defects in the sensor itself. This kind of fault was simulated by changing the measured bending moment signal to become constant after the fault occured.
@@
    Finally, it was decided to use real-life data collected from a blade undergoing a failure. For this reason, Vestas Wind Systems A/S provided some data collected with the turbine experiencing a fault of unknown origin, representing a real-life use case of a condition monitoring algorithm.
@@ 3 Methods and theoretical overview
Once the model is trained, the testing Stage is performed online: new data are collected from the wind turbine, into an unknown state, and then reconstructed by the trained model. The fault detection algorithm is obtained by computing a suitable statistic to compare the data reconstructed by the trained model and the real data measured by the sensors (usually by means of Square Prediction Error or Mean Average Error), and by comparing the result with the fault-free case obtained in the training stage by means of a Generalized likelihood ratio (GLR) test. If the two statistics are similar, then the GLR test assumes that the new data are from a healthy plant, while if they are showing strong differences, it detects a fault. Two approaches are proposed (PCA and Autoencoders), differing both in the model structure and in the training method.
@@ 3.1 Introduction to PCA
PCA is an unsupervised, linear method for dimensional reduction in large multivariate datasets. It transforms a normalized data matrix $\boldsymbol{X}\in\mathbb{R}^{n\times m}$ (with $n$ samples and $m$ features) into a lower-dimensional space by applying singular value decomposition (SVD) to its covariance matrix:
@@
$SPE(k)$ $=\sum_{i=1}^{m\omega}\left(\hat{X}_{d}^{(i)}(k)-X_{d}^{(i)}(k)\right)^{2}$ (4)
$=\boldsymbol{X_{d}}(k)(\boldsymbol{I}-\boldsymbol{\hat{P}\hat{P}}^{T})(\boldsymbol{I}-\boldsymbol{\hat{P}\hat{P}}^{T})^{T}\boldsymbol{X_{d}}(k)^{T}$
@@ 5 Experimental results
Both the PCA and the AE models are tested for the faulty scenarios presented in Section 2.1. The definition of the experiments for the third category is straightforward, thanks to the availability of faulty data. Two recordings are available, namely Real-life scenario 1 and Real-life scenario 2, with different evolution characteristics. On the other hand, the first two categories require the corruption of (unused) healthy data to mimic the faulty scenarios described in Section 2.1. Each of the first two scenarios (bias and stuck) are applied to both the edge and the flap bending moment. An overview of the experiment setting for each faulty scenario is presented in Table 2, where $M_{og}(t)$ represents the blade root bending moment measured on the original healthy turbine at time $t$, and $t_{f}$ the time instant where the fault occurs.
@@
Real-life scenario 1: catastrophic fault Sudden jump in the flap
bending moment signal
Real-life scenario 2: slower-forming Exponential increase in
the flap bending moment
Starting from the real-life scenarios, the reconstruction errors and the results of applying the GLR are presented in Figure 7. For the two scenarios under analysis, the faults are detected by both the PCA and AE models. This is observed when applying a threshold to the filtered reconstruction errors (SPE and MAE for the two methods, respectively) as well as when using the GLR, where the signal separation becomes more distinct.
The drawback of using real faulty data, is that the experiment is not repeatable. As a consequence, the statistical confidence of the methods can’t be evaluated.
@@
The same models are applied to the (new) healthy data corrupted with sensor faults, representing Scenarios 1 and 2. The faults are injected in each operating region, since the PCA requires to use a different model for each of them. The same testing scheme is applied to the AE, to test its capability of capturing the global nonlinear model. Because of the improved detection performance obtained with the GLR (thanks to its signal separation capability), the direct threshold methods results are omitted in the qualitative results presented in Figure 8. The capability of the AE of learning the nonlinear features of the model is reflected in a better performance over all the possible faults. In particular, the PCA is in general not able to clearly detect the fault associated with the Edge-bending-moment sensor, while it is still able to detect faults in the flap-bending moment sensor by leveraging the correct model according to the operating region the wind turbine is working at. Even under such conditions, the AE is better performing by ensuring strong detectability and no false alarms in the tests that have been performed.
@@
5.1 Detection statistical properties
In order to evaluate the performance of the proposed methods from a quantitative point of view, the possibility of manually corrupting healthy data in the first two faulty scenarios allows to have repeated experiments for each of them, and hence to derive some statistical properties for the defined architectures.
The metrics which is selected is the classic one from fault diagnosis, involving the probability of false alarms and the probability of detection [25]. The first refers to the amount of cases in which the threshold is surpassed before the fault is occurring, the latter to the number of crossing under fault, further distinguishing the two cases of strong and weak detection, where the signals is permanently surpassing the threshold or not, respectively. The total number of detection (without distinguishing between strong and weak) is calculated as well.
The same fault is injected under the same conditions for around 100 times, in order to obtain a precision in the probability percentages close to the unit. In the case of the PCA, where different operating regions are present, the total number of simualtions is set to be around 100 as well, with the number of simulations per operating regions selected according to data availability. Every single simulation lasts 900s, and the fault is injected at 300s.
The results are presented in Figure 9, for both PCA and AE and for both the direct threshold method and the GLR test.
Figure 9: Statistical properties of detection for PCA and AE models.
As it was noted from Figure 8, both models allow to always detect faults in the flap bending moment sensor. Missed detections are only present with the direct threshold for both of them. The case of the edge bending moment also confirms the outcome pf the previous analysis, with the PCA model showing a high percentage of missed detections with both detection methods (SPE and GLR).
The false alarm rate is generally higher for the AEs and when applying the GLR to the solution. The threshold may be varied to lower the value, at the cost of risking deterioration of the detection performance.
While the AE generally perform better at detection than the PCA, this is not verified for the case where the stuck flap bending moment, where the AE experiences a higher number of weak detections. However, this may be associated to the transient phase, after which the change is strongly detected.
@@ 6 Conclusions
The lack of investigations focusing on detecting blade faults in wind turbine lead to the definition of a condition monitoring framework based on learning based-models, trained to reconstruct a set of measurements under healthy conditions and showing discrepancies otherwise. Such architecture allows to detect abnormal behaviour in the blades for condition monitoring purposes.
[…]
Future work involves optimizing the algorithm for real-time implementation and testing different model architectures to improve detection performance, and being able to isolate the root causes.